OpenStudy (gabylovesyou):

Multiply x to the 3 fourths power times x to the 1 fifth power.

OpenStudy (gabylovesyou):

\[x^{\frac{ 3 }{ 4 }} * x^{\frac{ 1 }{ 5 }}\]

OpenStudy (anonymous):

Need help

OpenStudy (gabylovesyou):

yes please :) @lala098765

OpenStudy (anonymous):

All u need to do is mutlply them so x3/4 times x1\5

OpenStudy (gabylovesyou):

yes... how do i do that cx

Nnesha (nnesha):

when you multiply same variables you should add their exponent for example \[\huge\rm x^2 \times x^3 = x^{2+3}\]